OpenStudy (anonymous):

Solve the inequality. Express your answer in interval form. x^2 + 2x − 3 ≥ 0

OpenStudy (souvik):

factor left hand side

OpenStudy (anonymous):

x= -3 x=1

OpenStudy (anonymous):

(−infinity, −3] union [1, infinity) ?

OpenStudy (souvik):

you got..\[(x-1)(x+3)\ge 0\] right?

OpenStudy (anonymous):

yeah

OpenStudy (souvik):

it is possible when (x-1) and (x+3) must have same sign.....that means they must be positive both or negative both....agree?

OpenStudy (anonymous):

how are they both negative or positive its -3 on the inequality

OpenStudy (souvik):

let....\[x*y \ge0\] if we put x=5 and y=-5....does it satisfy?

OpenStudy (souvik):

it does not satisfy...if x and y have similar sign then it satisfy.... so \[(x-1)\ge 0\] or \(x\ge 1\) and \((x+3)\ge 0\) or \(x\ge -3\)